# BBP Phase Transition for a Doubly Sparse Deformed Model

### 05 Mar 2026 v1

@@ 1.1 Overview
A common setting alteration, and the model deviation most pertinent to this present work, considers sparse PCA where sparsity parameters, such as independent Bernoulli random variables, are applied entrywise to the spike vectors. In an application sense, this sparsity provides practical benefits in a high-dimensional setting, often improving interpretability, increasing efficiency, and reducing model complexity [ZHT06, JL09]. Under certain model assumptions, the general BBP threshold described in Informal Result 1 captures sparse PCA as the theorem assumes only a unit norm condition for the spike $v$; however, when considering alterations to the model’s parameters, Informal Result 1 does begin to fail. For instance, [BN11] covers both the eigenvector and eigenvalue result for multiple sparse spikes, but their analysis requires either the spike or noise to be orthogonally invariant. Seeing as orthogonal invariance is broken with the introduction of sparsity in the spike, this requires the noise matrix $W$ to be orthogonally invariant. Furthermore, their model requires the spike vectors to be orthogonal, thus limiting the sampling procedure and sparsity regimes for the spike. Methods for sparse PCA below the spectral threshold $\theta_{i}<1$ have been studied extensively as well (see Section 5.2). With the landscape of sparse PCA (under certain prior assumptions) being covered by BBP, we naturally consider the possibility of introducing sparsity across all aspects of the model, which brings us to the central question of this work:
@@ 1.2 Doubly Sparse Literature Review
The results of this present work are timely and could not have been produced if not for a series of recent contributions to the field of random matrix theory. First, a major ingredient involves the control of the operator norm for the sparse Wigner matrix. The study of extremal eigenvalues of sparse matrices is a field that has recently grown significantly; however, these works either consider cases with optimal sparsity ($q\gtrsim\frac{\log n}{n}$) but bounded random variable entries [TY21, ADK22] or sub-Gaussian entries but with sub-optimal sparsity ($q\gtrsim\frac{(\log n)^{4}}{n}$) [BBv23]. [AB26] provides the essential ingredient needed to show the top eigenvalue of our sub-Gaussian, sparse ($q\gg\frac{\log n}{n}$) Wigner matrix is bounded with high probability. Further, we require a local law result for our sparse Wigner matrix, which has been thoroughly explored over recent years [EKY+12, TVW13, EKY+13, DZ19, HK21]; we adapt the local law of [HWZ26] for own purposes as their result covers, to the best of our knowledge, the most optimal sparsity regime ($q\gg\frac{\log n}{n}$).
@@
We remark that, while our contributions and consideration of spectrally distinguishing and recovering a sparse spike from sparse noise are novel, the notion of a sparse noisy observation has been around for a while. Indeed, Robust PCA explores the problem of separating a low-rank structure from sparse noise. [CLM+11] provides a convex optimization procedure for separating these two entities, though the only randomness considered is the location of the sparsity in the mask. Their work gives an extensive background on the many applications of Robust PCA, including video surveillance, facial recognition, ranking and collaborative filtering, and more, showing how sparse noise grows increasingly prevalent in our large-dimensional modern age with instances of occlusion, tampering, and sensor failure. These aforementioned applications naturally extend to the setting of our model, where the spike also has a sparse structure. Moreover, doubly sparse applications have also been studied across different fields. [ZM18, SZM+19] provide an electrical engineering approach where they view the problem as a non-convex, NP-hard optimization problem on rectangular matrices with undefined sparsity to solve applications involving impulsive noise (image, audio, video denoising). More recently, [HAN25] looks at a comparable doubly sparse structure, but for non-Hermitian matrices, thus changing the landscape of the problem and relying on a completely different set of random matrix theory tools. [ASV25] examines, heuristically and empirically, a regime most similar to our setting, but instead considers a dense spike and a sparse, Poisson noise structure, which in turn produces an eigenvalue distribution with non-informative spikes. In their introduction, they highlight the recent interest in sparse noise and even go on to mention a subset of our model (i.e., sparse, sub-Gaussian noise) as a non-trivial direction of interest. However, in its current state, their model cannot capture our model parameters and setting even in denser regimes of sparsity due to their model construction procedure.
@@ 3 Main Results
where $p$ and $q$ are functions of $n,$ $r$ represents the number of the spikes, and $c_{1},C_{2},C_{3},C_{8},K$, and $D$ are positive constants defined in Theorems 4 and 7. Note $\bm{f}_{r}(n)\to 0$ as $n\to\infty$ under Assumption 5.
@@
in probability. More specifically, for any $\theta_{i}>1$ and $\gamma>0$, there exist some positive constants $c_{1},C_{2},C_{3},C_{5},C_{7},C_{8}$, and $D>0$ such that
@@
Remark.
As a concurrent work, [HLN26] studies the extreme eigenvalues of an $n\times n$ non‑Hermitian sparse i.i.d. matrix with sparsity $q$ and a deterministic finite‑rank additive deformation; they show that eigenvalue outliers outside the unit disk asymptotically match the eigenvalues of finite‑rank additive deformation if $qn\to\infty$. In contrast, our paper is Hermitian/symmetric (sparse Wigner) and the deformation is a set of random sparse spike vectors; Theorem 4 proves a BBP‑type transition with explicit outlier location $\theta+1/\theta$. Moreover, in the rank‑one case [HLN26] also controls the right‑eigenvector alignment similar to Theorem 7 but under an additional sparsity condition $qn\gg\log^{9}n$ while our Theorem 7 only requires $qn\gg\log(n)$.
@@ 5.1 Extensions Within BBP
Additionally, one may consider expanding the regime of sparsity for $q$ to the critical and sub-critical regimes. As discussed in Section 1.2, shifting to the assumption of bounded entries in the noise matrix allows for critical sparsity ($q\gtrsim\frac{\log n}{n})$ for our result of no spectral outliers, but we would still need a local law result in the style of Lemma 11 for critical sparsity. Further, we believe no BBP style result could be produced in the sub-critical regime as there would be no way to reliably distinguish between outliers generated by the spikes or spurious outliers from the noise. Thus, for $q\lesssim\frac{\log n}{n}$, a non-spectral approach would be required.
@@ 6.1 No Spectral Outliers for Sparse Wigner Matrix
Note that the following lemma holds by taking $\epsilon\to 0$.
Corollary 10.
Consider a sequence of matrices $\frac{1}{\sqrt{nq}}W_{n}\odot A_{n}$. Under the same assumptions of Lemma 9, we have
$\lambda_{1}(\frac{1}{\sqrt{nq}}W\odot A)\leq 2~,$
with high probability.
@@ 6.2 Local Law for Sparse Wigner Matrix
Proof.  Seeing as $W$ has sub-Gaussian entries, we use [HWZ26, Theorem 12] to choose a $D>0,\epsilon_{0}>0,$ and $\delta:=\delta(n)\to 0$. Then, there exist $n_{0}=n(D,\delta,K_{2},\epsilon_{0})$ defined in their work and constant $C_{1}=K_{2}^{4}*8^{2}*48^{2}*12^{6}\cdot\exp(2(D+7)+10+4/e)\cdot\max\{1,(\frac{5}{3}\delta)^{4}\}=K_{2}^{4}*8^{2}*48^{2}*12^{6}\exp(2(D+7)+10+4/e)$ for large enough $n$ as $\delta\to 0$ such that $q\geq C_{1}\frac{\log n}{\delta^{4}n}$ implies
$\mathbb{P}(\max_{1\leq i\leq n}|R_{n}(z)_{ii}-m(z)|\leq\delta)\geq 1-n^{-D}~,$
for
$z=E+i\eta\in S_{0}=\{E+i\eta:|E|\leq 10,n^{-1+\epsilon_{0}}\leq\eta\leq 1\}~,$
where $\epsilon_{0}\in(0,1)$ for $n\geq n_{0}$. In the context of their paper, we note $\beta=1/2$ as $W$ has sub-Gaussian (2-subexponential) entries, giving us the lower bound for $\eta$ in $S_{0}$.
@@ 6.3 Convergence of Top Eigenvalues
Proof.  Assume, for sake of contradiction, that there exists some fixed $i\geq 1$ such that $\lambda_{i}(\frac{1}{\sqrt{nq}}W\odot A)\not\xrightarrow[]{\mathbb{P}}2$. By Corollary 10, we know $\lambda_{1}(\frac{1}{\sqrt{nq}}W\odot A)\leq 2$ with high probability. Then, there exist $\bar{\epsilon},\bar{\delta}>0$ such that $\mathbb{P}(\lambda_{i}(\frac{1}{\sqrt{nq}}W\odot A)\not\in(2-\bar{\epsilon},2]:=I)>\bar{\delta}$. Thus, $\mathcal{N}_{I}\leq i-1$.
Observe the negation of Corollary 12: there exists a $D>0,\delta\in(0,1)$, or interval $I$ satisfying $|I|\geq C_{1}\frac{\log n}{\delta^{4}n}$ such that
@@
We now show the established interval $I:=(2-\bar{\epsilon},2]$ satisfies the conditions of this statement, thus reaching a contradiction. Let $C_{4}:=\int_{I}f_{\rm sc}(x)dx$.
Let $D>0$ be large enough such that $\bar{\delta}\geq n^{-D}$. Choose $\delta\leq\frac{C_{4}n-(i-1)}{n\bar{\epsilon}}\leq C_{4}/\bar{\epsilon}$ where the latter inequality holds for large enough $n$. Observe $|I|=\bar{\epsilon}\geq C_{1}\frac{(\bar{\epsilon})^{4}\log n}{C_{4}^{4}n}$ for large enough n.
@@
$\mathbb{P}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\mathcal{N}_{I}-n\int_{I}f_{\rm sc}(x)dx}}\right|\geq\delta n|I|}}\right)=$~ $\mathbb{P}\mathopen{}\mathclose{{\left(C_{4}n-\mathcal{N}_{I}\geq\delta n\bar{\epsilon}}}\right)$
$\geq$~ $\mathbb{P}\mathopen{}\mathclose{{\left(C_{4}n-\mathcal{N}_{I}\geq\frac{C_{4}n-(i-1)}{n\bar{\epsilon}}n\bar{\epsilon}}}\right)$
@@
$\geq$~ $\mathbb{P}\mathopen{}\mathclose{{\left(\lambda_{i}(\frac{1}{\sqrt{nq}}W\odot A)<(2-\bar{\epsilon},2]}}\right)$
@@ 7.2 Recovery Proof
$\mathopen{}\mathclose{{\left|\theta_{1}^{2}\langle u_{1},\frac{1}{\sqrt{np}}v_{1}\rangle^{2}-\frac{1}{m^{\prime}(z_{1})}}}\right|=$~ $\frac{1}{\sqrt{m^{\prime}(z_{1})}}\mathopen{}\mathclose{{\left|\theta_{1}^{2}m^{\prime}(z_{1})\langle u_{1},\frac{1}{\sqrt{np}}v_{1}\rangle^{2}+\theta_{1}^{2}\mathcal{M}_{1}\langle u_{1},\frac{1}{\sqrt{np}}v_{1}\rangle^{2}-\theta_{1}^{2}\mathcal{M}_{1}\langle u_{1},\frac{1}{\sqrt{np}}v_{1}\rangle^{2}-1}}\right|$
$\leq$~ $\frac{1}{\sqrt{m^{\prime}(z_{1})}}\mathopen{}\mathclose{{\left|\|\frac{1}{\sqrt{np}}\tilde{\Theta}V^{T}u_{1}\|^{2}-\sum_{j=2}^{r}\mathopen{}\mathclose{{\left(\theta_{j}^{2}(m^{\prime}(z_{1})+\mathcal{M}_{j})\langle u_{1},\frac{1}{\sqrt{np}}v_{j}\rangle^{2}}}\right)-1}}\right|+O(1)|\mathcal{M}_{1}|$
$\leq$~ $\frac{1}{\sqrt{m^{\prime}(z_{1})}}\mathopen{}\mathclose{{\left|\|\frac{1}{\sqrt{np}}\tilde{\Theta}V^{T}u_{1}\|^{2}-1}}\right|+\frac{r-1}{\sqrt{m^{\prime}(z_{1})}}\mathopen{}\mathclose{{\left(m^{\prime}(z_{1})+\mathcal{M}_{\max}}}\right)\cdot O(\Lambda^{2})+O(1)|\mathcal{M}_{1}|$
@@ Appendix A Appendix
A.1 Support Size Concentration of Spike Vectors
@@
A.2 Norm Concentration of Sparse Spike
@@
A.3 Bounding The Vanishing Terms
@@
The dominating term depends on the value of $\gamma>0$, the sparsity of $p$, and the sparsity parameter $\tau$. Thus, for values of $n$ large enough, we see there is a constant $C_{5}>0$ such that
@@ A.4 Non-distinct Signals
$\mathopen{}\mathclose{{\left|\theta_{1}^{2}\sum_{j=1}^{r_{0}}\langle u_{1},\frac{1}{\sqrt{np}}v_{j}\rangle^{2}-\frac{1}{m^{\prime}(z_{1})}}}\right|=$~ $\frac{1}{\sqrt{m^{\prime}(z_{1})}}\mathopen{}\mathclose{{\left|\theta_{1}^{2}\sum_{j=1}^{r_{0}}(m^{\prime}(z_{1})+\mathcal{M}_{j})\langle u_{1},\frac{1}{\sqrt{np}}v_{j}\rangle^{2}-\theta_{1}^{2}\sum_{j=1}^{r_{0}}\mathcal{M}_{j}\langle u_{1},\frac{1}{\sqrt{np}}v_{j}\rangle^{2}-1}}\right|$
@@
A.5 Varying Spike Sparsities

### 12 Mar 2026 v2

@@ 1.1 Overview
A common setting alteration, and the model deviation most pertinent to this present work, considers sparse PCA where sparsity parameters, such as independent Bernoulli random variables, are applied entrywise to the spike vectors. In an application sense, this sparsity provides practical benefits in a high-dimensional setting, often improving interpretability, increasing efficiency, and reducing model complexity [ZHT06, JL09]. Under certain model assumptions, the general BBP threshold described in Informal Result 1 captures sparse PCA as the theorem assumes only a unit norm condition for the spike $v$; however, when considering alterations to the model’s parameters, Informal Result 1 does begin to fail. For instance, [BN11] covers both the eigenvector and eigenvalue result for multiple sparse spikes, but their analysis requires either the spike or noise to be orthogonally invariant. Seeing as orthogonal invariance is broken with the introduction of sparsity in the spike, this requires the noise matrix $W$ to be orthogonally invariant. Methods for sparse PCA below the spectral threshold $\theta_{i}<1$ have been studied extensively as well (see Section 5.2). With the landscape of sparse PCA (under certain prior assumptions) being covered by BBP, we naturally consider the possibility of introducing sparsity across all aspects of the model, which brings us to the central question of this work:
@@ 1.2 Doubly Sparse Literature Review
The results of this present work are timely and could not have been produced if not for a series of recent contributions to the field of random matrix theory. First, a major ingredient involves the control of the operator norm for the sparse Wigner matrix. The study of extremal eigenvalues of sparse matrices is a field that has recently grown significantly; however, these works either consider cases with optimal sparsity ($q\gtrsim\frac{\log n}{n}$) but bounded random variable entries [TY21, ADK22] or sub-Gaussian entries but with sub-optimal sparsity ($q\gtrsim\frac{(\log n)^{4}}{n}$) [BBv23]. [AB26] provides the essential ingredient needed to show the top eigenvalue of our sub-Gaussian, sparse ($q\gg\frac{\log n}{n}$) Wigner matrix is bounded with high probability. Further, we require a local law result for our sparse Wigner matrix, which has been thoroughly explored over recent years [EKY+12, TVW13, EKY+13, DZ19, HK21]; we adapt the local law of [HWZ26] for own purposes as their result covers, to the best of our knowledge, the most optimal $\frac{\text{polylog}(n)}{n}$ regime.
@@
We remark that, while our contributions and consideration of spectrally distinguishing and recovering a sparse spike from sparse noise are novel, the notion of a sparse noisy observation has been around for a while. Indeed, Robust PCA explores the problem of separating a low-rank structure from sparse noise. [CLM+11] provides a convex optimization procedure for separating these two entities, though the only randomness considered is the location of the sparsity in the mask. Their work gives an extensive background on the many applications of Robust PCA, including video surveillance, facial recognition, ranking and collaborative filtering, and more, showing how sparse noise grows increasingly prevalent in our large-dimensional modern age with instances of occlusion, tampering, and sensor failure. These aforementioned applications naturally extend to the setting of our model, where the spike also has a sparse structure. Moreover, doubly sparse applications have also been studied across different fields. [ZM18, SZM+19] provide an electrical engineering approach where they view the problem as a non-convex, NP-hard optimization problem on rectangular matrices with undefined sparsity to solve applications involving impulsive noise (image, audio, video denoising). More recently, [HAN25] looks at a comparable doubly sparse structure, but for non-Hermitian matrices, thus changing the landscape of the problem and relying on a completely different set of random matrix theory tools. Similarly, [HLN26] studies the extreme eigenvalues of an $n\times n$ non‑Hermitian sparse i.i.d. matrix with sparsity $q$ and a deterministic finite‑rank additive deformation; they show that eigenvalue outliers outside the unit disk asymptotically match the eigenvalues of finite‑rank additive deformation if $qn\to\infty$. Moreover, in the rank‑one case [HLN26] provides control over the right‑eigenvector alignment similar to our Theorem 7 but under an additional sparsity condition $qn\gg\log^{9}n$ while we only require $qn\gg\log(n)$. [ASV25] examines, heuristically and empirically, a regime most similar to our setting, but instead considers a dense spike and a sparse, Poisson noise structure, which in turn produces an eigenvalue distribution with non-informative spikes. In their introduction, they highlight the recent interest in sparse noise and even go on to mention a subset of our model (i.e., sparse, sub-Gaussian noise) as a non-trivial direction of interest. However, in its current state, their model cannot capture our model parameters and setting even in denser regimes of sparsity due to their model construction procedure.
@@ 3 Main Results
where $p$ and $q$ are functions of $n,$ $r$ represents the number of spikes, and $c_{1},C_{2},C_{3},C_{8},K$, and $D$ are positive constants defined in Theorems 4 and 7. Note $\bm{f}_{r}(n)\to 0$ as $n\to\infty$ under Assumption 5.
@@
in probability. More specifically, for any $\theta_{i}>1$ and $\gamma>0$, there exist some constants $c_{1},C_{2},C_{3},C_{5},C_{7},C_{8}$, and $D>0$ such that
@@ 5.1 Extensions Within BBP
Additionally, one may consider expanding the regime of sparsity for $q$ to the critical and sub-critical regimes. As discussed in Section 1.2, shifting to the assumption of bounded entries in the noise matrix allows for critical sparsity ($q\gtrsim\frac{\log n}{n})$ for our result of no spectral outliers, but we would still need a local law result in the style of Lemma 10 for critical sparsity. Further, we believe no BBP style result could be produced in the sub-optimal regime as there would be no way to reliably distinguish between outliers generated by the spikes or spurious outliers from the noise. Thus, for $q\lesssim\frac{\log n}{n}$, a non-spectral approach would be required.
@@ 6.2 Local Law for Sparse Wigner Matrix
Proof.  Seeing as $W$ has sub-Gaussian entries, we use [HWZ26, Theorem 12] to choose a $D>0,\epsilon_{0}>0,$ and $\delta:=\delta(n)\to 0.$ Then, there exist $n_{0}=n(D,\delta,K_{2},\epsilon_{0})$ defined in their work and constant $C_{1}=K_{2}^{4}*8^{2}*48^{2}*12^{6}\cdot\exp(2(D+7)+10+4/e)\cdot\max\{1,(\frac{5}{3}\delta)^{4}\}=K_{2}^{4}*8^{2}*48^{2}*12^{6}\exp(2(D+7)+10+4/e)$ for large enough $n$ (since $\delta\to 0$) such that $q\geq C_{1}\frac{\log n}{\delta^{4}n}$ implies
$\mathbb{P}(\max_{1\leq i\leq n}|R_{n}(z)_{ii}-m(z)|\leq\delta)\geq 1-n^{-D}~,\text{ for }z=E+i\eta\in S_{0}=\{E+i\eta:|E|\leq 10,n^{-1+\epsilon_{0}}\leq\eta\leq 1\}~,$
where [^7] In the context of their paper, we note $\beta=1/2$ as $W$ has sub-Gaussian (2-subexponential) entries, giving us the lower bound for $\eta$ in $S_{0}$. $\epsilon_{0}\in(0,1)$ for $n\geq n_{0}$.
@@ 6.3 Convergence of Top Eigenvalues
Proof.  Assume, for sake of contradiction, that there exists some fixed $i\geq 1$ such that $\lambda_{i}(\frac{1}{\sqrt{nq}}W\odot A)\not\xrightarrow[]{\mathbb{P}}2$. Then, there exist $\bar{\epsilon},\bar{\delta}>0$ such that $\mathbb{P}(\lambda_{i}(\frac{1}{\sqrt{nq}}W\odot A)\not\in(2-\bar{\epsilon},2+\bar{\epsilon}):=I)>\bar{\delta}$. Thus, $\mathcal{N}_{I}\leq i-1$.
Observe the negation of Corollary 11: there exist a $D>0,\delta\in(0,1)$, or interval $I$ satisfying $|I|\geq C_{1}\frac{\log n}{\delta^{4}n}$ such that
@@
We now show the established interval $I:=(2-\bar{\epsilon},2+\bar{\epsilon})$ satisfies the conditions of this statement, thus reaching a contradiction. Let $C_{4}:=\int_{I}f_{\rm sc}(x)dx$.
Let $D>0$ be large enough such that $\bar{\delta}\geq n^{-D}$. Choose $\delta\leq\frac{C_{4}n-(i-1)}{2n\bar{\epsilon}}\leq C_{4}/(2\bar{\epsilon})$ where the latter inequality holds for large enough $n$. Observe $|I|=2\bar{\epsilon}\geq C_{1}{\log n}/({\delta^{4}n})$ for large enough n.
@@
$\mathbb{P}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\mathcal{N}_{I}-n\int_{I}f_{\rm sc}(x)dx}}\right|\geq\delta n|I|}}\right)=$~ $\mathbb{P}\mathopen{}\mathclose{{\left(C_{4}n-\mathcal{N}_{I}\geq 2\delta n\bar{\epsilon}}}\right)$
$\geq$~ $\mathbb{P}\mathopen{}\mathclose{{\left(C_{4}n-\mathcal{N}_{I}\geq 2\frac{C_{4}n-(i-1)}{2n\bar{\epsilon}}n\bar{\epsilon}}}\right)$
@@
$\geq$~ $\mathbb{P}\mathopen{}\mathclose{{\left(\lambda_{i}(\frac{1}{\sqrt{nq}}W\odot A)\not\in(2-\bar{\epsilon},2+\bar{\epsilon})}}\right)$
@@ 7.2 Recovery Proof
$\mathopen{}\mathclose{{\left|\theta_{1}^{2}\langle u_{1},\frac{1}{\sqrt{np}}v_{1}\rangle^{2}-\frac{1}{m^{\prime}(z_{1})}}}\right|=$~ $\frac{1}{m^{\prime}(z_{1})}\mathopen{}\mathclose{{\left|\theta_{1}^{2}m^{\prime}(z_{1})\langle u_{1},\frac{1}{\sqrt{np}}v_{1}\rangle^{2}+\theta_{1}^{2}\mathcal{M}_{1}\langle u_{1},\frac{1}{\sqrt{np}}v_{1}\rangle^{2}-\theta_{1}^{2}\mathcal{M}_{1}\langle u_{1},\frac{1}{\sqrt{np}}v_{1}\rangle^{2}-1}}\right|$
$\leq$~ $\frac{1}{m^{\prime}(z_{1})}\mathopen{}\mathclose{{\left|\|\frac{1}{\sqrt{np}}\tilde{\Theta}V^{T}u_{1}\|^{2}-\sum_{j=2}^{r}\mathopen{}\mathclose{{\left(\theta_{j}^{2}(m^{\prime}(z_{1})+\mathcal{M}_{j})\langle u_{1},\frac{1}{\sqrt{np}}v_{j}\rangle^{2}}}\right)-1}}\right|+O(1)|\mathcal{M}_{1}|$
$\leq$~ $\frac{1}{m^{\prime}(z_{1})}\mathopen{}\mathclose{{\left|\|\frac{1}{\sqrt{np}}\tilde{\Theta}V^{T}u_{1}\|^{2}-1}}\right|+\frac{r-1}{m^{\prime}(z_{1})}\mathopen{}\mathclose{{\left(m^{\prime}(z_{1})+\mathcal{M}_{\max}}}\right)\cdot O(\Lambda^{2})+O(1)|\mathcal{M}_{1}|$
@@ A.1 High Probability Conditioning
A.2 Support Size Concentration of Spike Vectors
@@
A.3 Norm Concentration of Sparse Spike
@@
A.4 Bounding The Vanishing Terms
@@
The dominating term depends on the value of $\gamma>0$, the sparsity of $p$, and the sparsity of $\tau$. Thus, for values of $n$ large enough, we see there is a constant $C_{5}>0$ such that
@@ A.5 Non-distinct Signals
$\mathopen{}\mathclose{{\left|\theta_{1}^{2}\sum_{j=1}^{r_{0}}\langle u_{1},\frac{1}{\sqrt{np}}v_{j}\rangle^{2}-\frac{1}{m^{\prime}(z_{1})}}}\right|=$~ $\frac{1}{m^{\prime}(z_{1})}\mathopen{}\mathclose{{\left|\theta_{1}^{2}\sum_{j=1}^{r_{0}}(m^{\prime}(z_{1})+\mathcal{M}_{j})\langle u_{1},\frac{1}{\sqrt{np}}v_{j}\rangle^{2}-\theta_{1}^{2}\sum_{j=1}^{r_{0}}\mathcal{M}_{j}\langle u_{1},\frac{1}{\sqrt{np}}v_{j}\rangle^{2}-1}}\right|$
@@
A.6 Varying Spike Sparsities